jowhite:

Make a table of the function. y = x + 1

AZ:

to make a table, just start using random numbers in place of 'x' so when x = 0, what is y? y = x + 1 y = 0 + 1 y = ?? and similarly, just find the y-value when x is equal to different numbers

jowhite:

im confused

AZ:

I'm assuming you don't have any answer choices otherwise, a table looks like this 

AZ:

you just choose any number you want for 'x' and then you find the number for 'y' and you write it down in the table

jowhite:

yea ik

AZ:

so for example, when x = 0 we get y = 0 + 1 y = 1 so we would write that in our table as 

jowhite:

Ok